Calcium cyanamide reacts with water to form calcium carbonate and ammonia. a.How many grams of H2O are needed to react with 75.0 g of CaCN2? b.How many grams of NH3 are produced from 5.24 g of CaCN2? c.How many grams of CaCO3 form if 155 g of H2O reacts?


1
Expert's answer
2021-09-27T06:33:21-0400

CaCN2(s)+3H2O(l)-->CaCO3(s)+2NH3(g)

Molar mass of CaCN2 = 92.11

75/92.11 = 0.814m

Mole ratio = 1:3

= 3×0.814 = 2.443

Molar mass of H2O = 18.02

2.443×18.02 = 44.02g